Circular No. 7620

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVA 2001bf IN MCG +04-42-22 G. M. Hurst, Basingstoke, England, reports the discovery of an ap-

parent supernova (mag 16.5) by M. Armstrong, Rolvenden, on an unfiltered CCD frame (limiting mag 18) taken on May 3.020 UT with a 0.30-m Schmidt-Cassegrain reflector in the course of the U.K. Nova/Supernova Patrol. Armstrong measured the new object’s position to be α = 18^{h}01^{m}33^{s}.99, δ = +26^{o}15′02′′.3 (equinox 2000.0), which is 5′′ east and 8′′ south of the center of MCG +04-42-22. An image taken by Armstrong on May 4.965 shows that SN 2001bf had brightened to mag 15.9. An image taken on May 5.001 by T. Boles, Coddenham, also confirms the presence of the new object. A master patrol image taken by Armstrong on 2000 June 8 showed nothing at this position (limiting mag 18.5), and nothing appears on Palomar Sky Survey images from 1991 May 14 (limiting red mag 20.8) and 1991 July 12 (limiting blue mag 22.5).

SUPERNOVA 2001bb IN IC 4319 Regarding his discovery (cf. IAUC 7614), R. Chassagne adds that a

prediscovery CCD image shows SN 2001bb at mag 16.5–17.0 on Apr. 27.8.

V445 PUPPIS A. C. Gilmore and P. M. Kilmartin report photometry obtained with

the 0.6-m f/16 Cassegrain telescope at the Mt. John Observatory, indicating that this novalike variable is fading after varying around mag 10 for the past two months: Apr. 12.41 UT, V = 9.80, U −B = −0.40, B−V = +0.56, V − R = +0.46, V − I = +1.03; 13.36, 9.97, −0.39, +0.56, +0.47, +1.05; 14.37, 10.13, −0.39, +0.55, +0.50, +1.07; 16.38, 10.21, −0.39, +0.54, +0.50, +1.04; 18.50, 10.33, −0.40, +0.53, +0.53, +1.09; 21.41, 10.64, −0.33, +0.56, +0.54, +1.13; 22.40, 10.61, −0.26, +0.59, +0.56, +1.17; 23.45, 10.74, −0.26, +0.62, +0.60, +1.22; 24.35, 10.86, −0.32, +0.60, +0.62, +1.27; 26.33, 11.04, −0.14, +0.62, +0.63, +1.29. The uncertainties are generally < 0.02 mag (but Apr. 24 observations affected by cloud).

COMET C/2001 A2 (LINEAR) Visual m_{1} estimates: May 1.46 UT, 5.9 (M. Mattiazzo, Wallaroo, S.

Australia, 7×50 binoculars); 3.49, 6.2 (A. Pearce, Nedlands, W. Australia, 20×80 binoculars); 4.38, 5.6 (D. A. J. Seargent, The Entrance, N.S.W., 10×50 binoculars); 5.25, 5.8 (M. Linnolt, Honolulu, HI, 10×50 binoculars).

2001 May 5  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2001 CBAT  Daniel W. E. Green